# Understanding and Mitigating Memorization in Generative Models via Sharpness of Probability Landscapes

### v3

@@ 4.1 Memorization: Sharpness in Probability Landscape
Figure 3: Left: Eigenvalue distribution of $H_{\theta}(\mathbf{x}_{t},c)$ across memorization categories in Stable Diffusion v1.4 at initial sampling step ($t=T-1$) with range clipped. (top) 30 prompts per category with identical initialization. (bottom) Fixed prompt set with three different initializations. Both plots reveal stronger memorization correlates with fewer non-negative eigenvalues. Right: Eigenvalue distribution of $H_{\theta}(\mathbf{x}_{t},c)$ across memorization categories in Stable Diffusion v1.4 at final sampling step ($t=1$). Generated images shown with original training counterparts (outlined in red). Eigenvalues are approximated via Arnoldi iteration (Arnoldi, 1951), details in Section A.2.
Figure 2: Left: Generated images for memorized (digit “9”) and non-memorized (digit “3”) samples. Right: Eigenvalue distributions for memorized (red) and non-memorized (blue) samples at initial (top) and final (bottom) sampling steps, revealing more and larger negative eigenvalues in memorized cases. Experimental details in Appendix C.
@@
To validate our approach on real data, we conduct experiments on MNIST by inducing memorization through repeated exposure to a single “9” image while maintaining all “3” images as a general class (Figure 2). The eigenvalue distributions at $t=1$ clearly differentiate memorized from non-memorized samples: memorized samples show consistently large negative eigenvalues indicating sharp peaks, while non-memorized samples exhibit positive eigenvalues, reflecting locally convex regions that allow sample variations. Notably, these clear distributional differences emerged even at the initial sampling step ($t=T-1$), confirming that sharpness-based memorization detection is effective from the very beginning of the generation process.
@@ 4.2 Score Norm as a Sharpness Measure
Under this Gaussian assumption at relevant sampling steps, the score norm $\|s_{\theta}(\mathbf{x}_{t})\|^{2}$ provides an unbiased estimate of the negative Hessian trace $-\textup{{{tr}}}(H_{\theta}(\mathbf{x}_{t}))$, offering an efficient measure of the sharpness of the probability landscape.
Figure 4: Empirical alignment in MNIST and Stable Diffusion between: (left) $-\textup{{{tr}}}\bigr{(}H_{\theta}(\mathbf{x}_{t},c)\bigl{)}$ and $\|s_{\theta}(\mathbf{x}_{t},c)\|^{2}$, and (right) $-\textup{{{tr}}}\bigr{(}H_{\theta}(\mathbf{x}_{t},c)^{3}\bigl{)}$ and $\|H_{\theta}(\mathbf{x}_{t},c)s_{\theta}(\mathbf{x}_{t},c)\|^{2}$.
Figure 4 empirically confirms that this approximation holds reliably across datasets, including MNIST and Stable Diffusion’s latent space. Surprisingly, this relationship persists even in the later stages of the diffusion process, suggesting that score norm can serve as a computationally efficient sharpness measure throughout generation. This perspective provides a theoretical foundation for interpreting sharpness in generative models through score norm based statistic, enabling efficient memorization detection and analysis without requiring costly Hessian eigenvalue decompositions.
@@ 4.3 Wen’s Metric as a Sharpness Measure
Figure 5 shows the eigenvalue disparities between conditional and unconditional Hessians across timesteps, revealing how conditioning shapes the probability distribution’s geometry. For memorized samples, the eigenvalue gap is notably large, showing that conditioning creates a more constrained probability landscape. At intermediate timesteps ($t=20$), the differences are subtle but noticeable, indicating early conditioning effects. Near the end ($t=1$), the eigenvalue gap widens substantially, demonstrating conditioning’s growing influence on the learned density. In contrast, non-memorized samples show minimal eigenvalue variations throughout, indicating little conditioning influence. These findings support our theoretical framework and confirm Wen’s metric effectively measures sharpness.
@@ 4.4 Upscaling Eigenvalue Statistics via Hessian
Table 1: AUC and TPR@1%FPR across detection strategies and sampling steps for Stable Diffusion (SD) v1.4 and v2.0. Here, $n$ denotes the number of generations per prompt, with results averaged over $n$. “Steps” indicates the stage along the diffusion sampling path, ranging from step $1\,(t=T-1)$ to step $50\,(t=1).$
@@
This relationship, empirically verified in  Figure 4, demonstrates the curvature-sensitive scaling effect of the Hessian score product. Building on this principle, we propose an enhanced version of Wen’s metric that improves early-stage sensitivity through second-order sharpness characterization:
@@
Compared to Wen’s metric in Lemma 4.2, this refinement substantially improves sensitivity by amplifying the difference in sharpness, thereby enabling more effective detection of memorization at earlier stages.
@@ 4.5 Detecting Memorization in Stable Diffusion
We report detection results at sampling steps 1, 5, and 50, but only include 50-step results for methods requiring full sampling or showing significant performance gains. Additional experimental details are provided in Section D.1.
@@ Results.
Table 1 demonstrates our metric’s strong performance on Stable Diffusion v1.4 and v2.0 using just a single sampling step. By upscaling curvature information via $H_{\theta}^{\Delta}(\mathbf{x}_{t})$, we significantly enhance Wen et al. (2024)’s metric. With merely four generations, we achieve an AUC of 0.998 and TPR@1%FPR of 0.982, matching Wen et al. (2024)’s performance using five steps and 16 generations. Similarly, in v2.0, our approach attains an AUC of 0.991 without full-step sampling, underscoring its effectiveness.
@@ 5 Sharpness Aware Memorization Mitigation
Figure 6: Left: Comparison of inference-time mitigation methods on SD v1.4 (top) and v2.0 (bottom), evaluated across five hyperparameter configurations per method. Lower SSCD scores indicate reduced memorization, while higher CLIP scores show better prompt-image alignment. Right: Qualitative comparison demonstrating SAIL’s effectiveness in preserving key image details (shown adjacent to the original image), whereas baseline methods exhibit quality degradation due to modified text conditioning. Images are generated using identical random seeds, with full prompts in Section D.2
@@ 5.1 Sharpness Aware Initialization Sampling
In Section 4, we observed that memorized samples exhibit a sharp conditional density, $p_{t}(\mathbf{x}_{t}|c)$, even at the very beginning of the generation process (i.e., at $t=T-1$; note that sampling proceeds in reverse order, starting from $t=T$). This is substantiated by the strong detection performance of both Wen’s metric and our metric at the initial sampling step, which quantifies the sharpness gap between $p_{t}(\mathbf{x}_{t}|c)$ and $p_{t}(\mathbf{x}_{t})$.
@@
We propose Sharpness-Aware Initialization for Latent Diffusion (SAIL), an inference-time mitigation strategy that optimizes initializations $\mathbf{x}_{T}$ by minimizing the sharpness gap at the starting step ($t=T-1$). SAIL identifies initial seeds on non-memorized sampling trajectories by selecting $\mathbf{x}_{T}$ from smoother regions while maintaining a reasonable density under the isotropic Gaussian prior. The objective function is defined as:
@@
To evaluate mitigation strategies, we use the same memorized prompt set employed in the detection experiments described in Section 4.5. However, since verifying mitigation effects requires access to training images, we exclude prompts whose corresponding training samples are unavailable. Further details are in Appendix D.
@@
For a fair comparison, all methods are evaluated using five distinct hyperparameter settings and optimized with the Adam optimizer at a learning rate of 0.05. For a detailed experimental settings, refer to Section D.2.
@@ Results.
Figure 6 (left) demonstrates that SAIL significantly improves both SSCD and CLIP metrics for Stable Diffusion v1.4 and v2.0. By optimizing the noise initialization $\mathbf{x}_{T}$ without altering model components like text embeddings or attention weights, SAIL effectively mitigates memorized content while preserving model behavior and user prompts, ensuring high-quality, non-memorized outputs.
The advantage of SAIL is evident in Figure 6 (right), where it generates images that faithfully preserve key prompt details, such as celebrity names and primary objects. In contrast, methods that modify text-conditional components often reduce the influence of those components during mitigation, leading to degraded alignment with the original prompt and potentially diminishing user utility. Additional qualitative results for algorithms are provided in Appendix E.
@@ A.1 Second-Order Score Function
[…]
@@ A.2 Numerical Eigenvalue Algorithm
[…]
@@
After orthogonalization, we compute $h_{k+1,k}=\|\mathbf{v}\|_{2}$. If $h_{k+1,k}$ is greater than a small threshold $\varepsilon$, we normalize $\mathbf{v}$ to obtain the next basis vector $\mathbf{q}_{k+1}=\frac{\mathbf{v}}{h_{k+1,k}}$. Otherwise, the iteration terminates. The eigenvalues of $\mathbf{H}_{m}$(Ritz values) approximate the $m$ eigenvalues of $\mathbf{A}$. For details on the computational process of Arnoldi iteration, Please refer to the algorithm pesudo code below. The Arnoldi iteration tends to find eigenvalues with larger absolute values first because components associated with these eigenvalues dominate within the Krylov subspace. If the input matrix is symmetric, Arnoldi iteration can be simplified to Lanczos iteration (Lanczos, 1950). However, since the Lanczos iteration is very sensitive to small numerical errors breaking the symmetry, we use the general version. The computational complexity of the algorithm is $O(m^{2}d)$ with space complexity $O(md)$, compared to $O(d^{3})$ with $O(d^{2})$ of exact derivation and eigendecomposition of Hessian. We calculate all eigenvalues for several samples for clear justification. But with just a few$(m\ll d)$ iterations, the difference between memorized samples and non-memorized samples reveals enough.
@@
[…]
@@ A Gaussian log-density has
$\log p(\mathbf{x})=\;-\tfrac{1}{2}(\mathbf{x}-\boldsymbol{\mu})^{\top}\!\boldsymbol{\Sigma}^{-1}(\mathbf{x}-\boldsymbol{\mu})+\text{const.},$. ,
@@
For a random vector $\mathbf{x}\sim p(\mathbf{x})$ with regularity conditions $\mathbb{E}\left[||s(\mathbf{x})||^{2}\right]<\infty$ and $\underset{\|x\|\to\infty}{\lim}p(\mathbf{x})s(\mathbf{x})=\mathbf{0}$,
@@ Appendix C Details of the Toy Experiments
This section provides additional details on the 2D and MNIST experiments discussed in Section 4.1. For both experiments, we use the DDPM (Ho et al., 2020) framework with the DDIM (Song et al., 2021a) sampler, employing 500 sampling steps. Additionally, to obtain a more accurate estimate of the Hessian (Jacobian of the score function), we utilize the second-order score matching loss proposed by Lu et al. (2022) during model training. Refer to Section A.1 for details.
@@
In Figure 2, for the non-memorized case, we sample 1,000 images and select the top 500 samples with the largest pairwise $\ell_{2}$ distances from training samples to highlight cases clearly deviating from memorization. For the memorized case, as all images collapse into a single image, we sample 500 outputs without comparing $\ell_{2}$ distances.
@@ Appendix D Details of the Stable Diffusion Experiments
This section describes the experimental setups for the Stable Diffusion experiments presented in Section 4.5 and Section 5. We provide a detailed overview of the configurations, including the specific prompts used and the implementation details of the baseline methods.
@@ D.1 Memorization Detection
  1. Tiled $\mathbf{\ell_{2}}$ distance: Building on the insight that memorized prompts produce similar generations regardless of their initializations, Carlini et al. (2023) propose examining generation density by analyzing multiple generated images for a given prompt using pairwise $\ell_{2}$ distances in pixel space. To address false positives from similar backgrounds, Carlini et al. (2023) divide images into non-overlapping $128\times 128$ tiles and compute the maximum $\ell_{2}$ distance between corresponding tiles. We adopt the identical setting for both Stable Diffusion v1.4 and v2.0. As the detection performance of this metric achieves the best after full sampling steps, we only report the complete 50-step results in Table 1.
@@
    where $\overline{a}_{i}^{l}$ is the average attention score in layer $l$. For detection experiments, we follow the implementation and hyperparameter settings of Ren et al. (2024). The detection performance differences between our results in Table 1 and those reported in Ren et al. (2024) can be attributed to different choices of non-memorized prompts. Specifically, our evaluation uses prompts collected from diverse sources, whereas Ren et al. (2024) utilizes GPT-4 generated prompts that share similar characteristics. For comprehensive experimental details, we refer readers to Ren et al. (2024).
@@
    Consequently, Wen’s metric can be defined as the norm of score differences as described in Section 4.3.
@@
    We note that the attention mask $\mathbf{m}$ is obtainable at the final sampling step ($t=1$). Therefore, to utilize BE as a detection metric, the model requires completion of all sampling steps. Consequently, in Table 1, we report experimental results using the complete 50-step diffusion process.

### v4

@@ 4.1 Memorization: Sharpness in Probability Landscape
Figure 3: Left: Eigenvalue distribution of $H_{\theta}(\mathbf{x}_{t},c)$ across memorization categories in Stable Diffusion v1.4 at initial sampling step (t=T1t=T-1) with range clipped. (top) 30 prompts per category with identical initialization. (bottom) Fixed prompt set with three different initializations. Both plots reveal stronger memorization correlates with fewer non-negative eigenvalues. Right: Eigenvalue distribution of $H_{\theta}(\mathbf{x}_{t},c)$ across memorization categories in Stable Diffusion v1.4 at final sampling step ($t=1$). Generated images shown with original training counterparts (outlined in red). Eigenvalues are approximated via Arnoldi iteration (Arnoldi, 1951), details in Section˜A.2.
Figure 2: Left: Generated images for memorized (digit “9”) and non-memorized (digit “3”) samples. Right: Eigenvalue distributions for memorized (red) and non-memorized (blue) samples at initial (top) and final (bottom) sampling steps, revealing more and larger negative eigenvalues in memorized cases. Experimental details in Appendix˜C.
@@
To validate our approach on real data, we conduct experiments on MNIST by inducing memorization through repeated exposure to a single “9” image while maintaining all “3” images as a general class (Figure 2). The eigenvalue distributions at $t=1$ clearly differentiate memorized from non-memorized samples: memorized samples show consistently large negative eigenvalues indicating sharp peaks, while non-memorized samples exhibit positive eigenvalues, reflecting locally convex regions that allow sample variations. Notably, these clear distributional differences emerged even at the initial sampling step (t=T1t=T-1), confirming that sharpness-based memorization detection is effective from the very beginning of the generation process.
@@ 4.2 Score Norm as a Sharpness Measure
Under this Gaussian assumption at relevant sampling steps, the score norm $\|s_{\theta}(\mathbf{x}_{t})\|^{2}$ provides an unbiased estimate of the negative Hessian trace $\textup{{{tr}}}(H_{\theta}(\mathbf{x}_{t}))$, offering an efficient measure of the sharpness of the probability landscape.
Figure 4: Empirical alignment in MNIST and Stable Diffusion between: (left) $\textup{{{tr}}}\bigr{(}H_{\theta}(\mathbf{x}_{t},c)\bigl{)}$ and $\|s_{\theta}(\mathbf{x}_{t},c)\|^{2}$, and (right) $\textup{{{tr}}}\bigr{(}H_{\theta}(\mathbf{x}_{t},c)^{3}\bigl{)}$ and $\|H_{\theta}(\mathbf{x}_{t},c)s_{\theta}(\mathbf{x}_{t},c)\|^{2}$.
Figure˜4 empirically confirms that this approximation holds reliably across datasets, including MNIST and Stable Diffusion’s latent space. Surprisingly, this relationship persists even in the later stages of the diffusion process, suggesting that score norm can serve as a computationally efficient sharpness measure throughout generation. This perspective provides a theoretical foundation for interpreting sharpness in generative models through score norm based statistic, enabling efficient memorization detection and analysis without requiring costly Hessian eigenvalue decompositions.
@@ 4.3 Wen’s Metric as a Sharpness Measure
Figure˜5 shows the eigenvalue disparities between conditional and unconditional Hessians across timesteps, revealing how conditioning shapes the probability distribution’s geometry. For memorized samples, the eigenvalue gap is notably large, showing that conditioning creates a more constrained probability landscape. At intermediate timesteps ($t=20$), the differences are subtle but noticeable, indicating early conditioning effects. Near the end ($t=1$), the eigenvalue gap widens substantially, demonstrating conditioning’s growing influence on the learned density. In contrast, non-memorized samples show minimal eigenvalue variations throughout, indicating little conditioning influence. These findings support our theoretical framework and confirm Wen’s metric effectively measures sharpness.
@@ 4.4 Upscaling Eigenvalue Statistics via Hessian
Table 1: AUC and TPR@1%FPR across detection strategies and sampling steps for Stable Diffusion (SD) v1.4 and v2.0. Here, $n$ denotes the number of generations per prompt, with results averaged over $n$. “Steps” indicates the stage along the diffusion sampling path, ranging from step 1(t=T1)1\,(t=T-1) to step 50(t=1).50\,(t=1).
@@
This relationship, empirically verified in  Figure˜4, demonstrates the curvature-sensitive scaling effect of the Hessian score product. Building on this principle, we propose an enhanced version of Wen’s metric that improves early-stage sensitivity through second-order sharpness characterization:
@@
Compared to Wen’s metric in Lemma˜4.2, this refinement substantially improves sensitivity by amplifying the difference in sharpness, thereby enabling more effective detection of memorization at earlier stages.
@@ 4.5 Detecting Memorization in Stable Diffusion
We report detection results at sampling steps 1, 5, and 50, but only include 50-step results for methods requiring full sampling or showing significant performance gains. Additional experimental details are provided in Section˜D.1.
@@ Results.
Table˜1 demonstrates our metric’s strong performance on Stable Diffusion v1.4 and v2.0 using just a single sampling step. By upscaling curvature information via $H_{\theta}^{\Delta}(\mathbf{x}_{t})$, we significantly enhance Wen et al. (2024)’s metric. With merely four generations, we achieve an AUC of 0.998 and TPR@1%FPR of 0.982, matching Wen et al. (2024)’s performance using five steps and 16 generations. Similarly, in v2.0, our approach attains an AUC of 0.991 without full-step sampling, underscoring its effectiveness.
@@ 5 Sharpness Aware Memorization Mitigation
Figure 6: Left: Comparison of inference-time mitigation methods on SD v1.4 (top) and v2.0 (bottom), evaluated across five hyperparameter configurations per method. Lower SSCD scores indicate reduced memorization, while higher CLIP scores show better prompt-image alignment. Right: Qualitative comparison demonstrating SAIL’s effectiveness in preserving key image details (shown adjacent to the original image), whereas baseline methods exhibit quality degradation due to modified text conditioning. Images are generated using identical random seeds, with full prompts in Section˜D.2
@@ 5.1 Sharpness Aware Initialization Sampling
In Section˜4, we observed that memorized samples exhibit a sharp conditional density, $p_{t}(\mathbf{x}_{t}|c)$, even at the very beginning of the generation process (i.e., at t=T1t=T-1; note that sampling proceeds in reverse order, starting from $t=T$). This is substantiated by the strong detection performance of both Wen’s metric and our metric at the initial sampling step, which quantifies the sharpness gap between $p_{t}(\mathbf{x}_{t}|c)$ and $p_{t}(\mathbf{x}_{t})$.
@@
We propose Sharpness-Aware Initialization for Latent Diffusion (SAIL), an inference-time mitigation strategy that optimizes initializations $\mathbf{x}_{T}$ by minimizing the sharpness gap at the starting step (t=T1t=T-1). SAIL identifies initial seeds on non-memorized sampling trajectories by selecting $\mathbf{x}_{T}$ from smoother regions while maintaining a reasonable density under the isotropic Gaussian prior. The objective function is defined as:
@@
To evaluate mitigation strategies, we use the same memorized prompt set employed in the detection experiments described in Section˜4.5. However, since verifying mitigation effects requires access to training images, we exclude prompts whose corresponding training samples are unavailable. Further details are in Appendix˜D.
@@
For a fair comparison, all methods are evaluated using five distinct hyperparameter settings and optimized with the Adam optimizer at a learning rate of 0.05. For a detailed experimental settings, refer to Section˜D.2.
@@ Results.
Figure˜6 (left) demonstrates that SAIL significantly improves both SSCD and CLIP metrics for Stable Diffusion v1.4 and v2.0. By optimizing the noise initialization $\mathbf{x}_{T}$ without altering model components like text embeddings or attention weights, SAIL effectively mitigates memorized content while preserving model behavior and user prompts, ensuring high-quality, non-memorized outputs.
The advantage of SAIL is evident in Figure˜6 (right), where it generates images that faithfully preserve key prompt details, such as celebrity names and primary objects. In contrast, methods that modify text-conditional components often reduce the influence of those components during mitigation, leading to degraded alignment with the original prompt and potentially diminishing user utility. Additional qualitative results for algorithms are provided in Appendix˜E.
@@ A.1 Second-Order Score Function
$\theta^{*}=\underset{\theta}{\arg\min}\,\mathbb{E}_{\mathbf{x}_{0},\boldsymbol{\epsilon}}\left[\frac{1}{\sigma_{t}^{4}}\left\|\sigma_{t}^{2}H_{\theta}(\mathbf{x}_{t})+\mathbf{I}-\ell_{1}\ell_{1}^{\top}\right\|_{F}^{2}\right],$
@@ A.2 Numerical Eigenvalue Algorithm
where $\mathbf{e}_{m}$ is the $m$-th canonical basis vector. Since we can compute $\mathbf{A}\mathbf{q}_{k}$ without forming $\mathbf{A}$ explicitly, using the function $\text{jvp\_func}(\mathbf{q}_{k})$, the Arnoldi iteration proceeds as follows. First, we normalize the starting vector $\mathbf{b}$ to obtain $\mathbf{q}_{1}=\frac{\mathbf{b}}{\|\mathbf{b}\|_{2}}$. Then, for each iteration $k=1$ to $m$, we compute:
@@
[…]
@@
0: Starting vector $\mathbf{b}\in\mathbb{R}^{d}$, number of iterations $m\leq d$, function $\text{jvp\_func}(\mathbf{v})$ that computes $\mathbf{A}\mathbf{v}$, threshold $\varepsilon$
@@ A Gaussian log-density has
$\log p(\mathbf{x})=\;-\tfrac{1}{2}(\mathbf{x}-\boldsymbol{\mu})^{\top}\!\boldsymbol{\Sigma}^{-1}(\mathbf{x}-\boldsymbol{\mu})+\text{const.},$
@@
[…]
@@ Appendix C Details of the Toy Experiments
This section provides additional details on the 2D and MNIST experiments discussed in Section˜4.1. For both experiments, we use the DDPM (Ho et al., 2020) framework with the DDIM (Song et al., 2021a) sampler, employing 500 sampling steps. Additionally, to obtain a more accurate estimate of the Hessian (Jacobian of the score function), we utilize the second-order score matching loss proposed by Lu et al. (2022) during model training. Refer to Section˜A.1 for details.
@@
In Figure˜2, for the non-memorized case, we sample 1,000 images and select the top 500 samples with the largest pairwise $\ell_{2}$ distances from training samples to highlight cases clearly deviating from memorization. For the memorized case, as all images collapse into a single image, we sample 500 outputs without comparing $\ell_{2}$ distances.
@@ Appendix D Details of the Stable Diffusion Experiments
This section describes the experimental setups for the Stable Diffusion experiments presented in Section˜4.5 and Section˜5. We provide a detailed overview of the configurations, including the specific prompts used and the implementation details of the baseline methods.
@@ D.1 Memorization Detection
  1. Tiled $\mathbf{\ell_{2}}$ distance: Building on the insight that memorized prompts produce similar generations regardless of their initializations, Carlini et al. (2023) propose examining generation density by analyzing multiple generated images for a given prompt using pairwise $\ell_{2}$ distances in pixel space. To address false positives from similar backgrounds, Carlini et al. (2023) divide images into non-overlapping $128\times 128$ tiles and compute the maximum $\ell_{2}$ distance between corresponding tiles. We adopt the identical setting for both Stable Diffusion v1.4 and v2.0. As the detection performance of this metric achieves the best after full sampling steps, we only report the complete 50-step results in Table˜1.
@@
    where $\overline{a}_{i}^{l}$ is the average attention score in layer $l$. For detection experiments, we follow the implementation and hyperparameter settings of Ren et al. (2024). The detection performance differences between our results in Table˜1 and those reported in Ren et al. (2024) can be attributed to different choices of non-memorized prompts. Specifically, our evaluation uses prompts collected from diverse sources, whereas Ren et al. (2024) utilizes GPT-4 generated prompts that share similar characteristics. For comprehensive experimental details, we refer readers to Ren et al. (2024).
@@
    Consequently, Wen’s metric can be defined as the norm of score differences as described in Section˜4.3.
@@
    We note that the attention mask $\mathbf{m}$ is obtainable at the final sampling step ($t=1$). Therefore, to utilize BE as a detection metric, the model requires completion of all sampling steps. Consequently, in Table˜1, we report experimental results using the complete 50-step diffusion process.